OpenStudy (anonymous):

Two rectangles are similar. Rectangle ABCD has a length of 5 in and a width of 10 in. Rectangle ABCD is dilated by a scale factor of 4 to form rectangle FGHJ. Answer these questions using the information provided: 1. How is the area affected by the dilation? 2. How is the perimeter affected by the dilation?

OpenStudy (raden):

to get the length and the width of rectangle of FGHJ, just multiply by 4 from the length and the width of rectangle ABCD. so, l = 4 * 5 = 20 w = 4 * 10 = 40 now, to get area use the formula : area = length * width = 20 * 40 = .... in

OpenStudy (anonymous):

Thanks; let me read your reply.

OpenStudy (anonymous):

It might take a few mins,ok?

OpenStudy (anonymous):

60 in,yes?

OpenStudy (raden):

for perimeter of rectangle FGHJ, use the formula : perimeter = 2(l+w) = 2(20+40) = 2 * 60 = ... in

OpenStudy (raden):

yes, that's right sorry, the unit for area must be in squared not in :)

OpenStudy (anonymous):

ok, thank u! so, how should I state my answer?

OpenStudy (anonymous):

You can call me '123', by the way.

OpenStudy (raden):

oppss.. my eyes still want to bed, @123 :) area = 20 * 40 = 80 in^2 not 60 (sorry) perimeter = 2(20+40) = 2(60) = 120 in

OpenStudy (anonymous):

lol. :D

OpenStudy (raden):

opss.. 20 * 40 = 800 not 80 my stupid --"

OpenStudy (anonymous):

nope. so, what should i write for number 1?

OpenStudy (raden):

800 in^2

OpenStudy (anonymous):

oK. But 1 is asking how the area was affected by the dilation.

OpenStudy (anonymous):

@AriPotta

OpenStudy (raden):

it can be : area = 16 times of the area of before do dilated

OpenStudy (anonymous):

Oh,ok. Let me write that down .

OpenStudy (anonymous):

or '16 times the original area'?

OpenStudy (raden):

yes

OpenStudy (aripotta):

4^2

OpenStudy (anonymous):

Ok, thx

OpenStudy (aripotta):

i told you that trick before, enchilada

OpenStudy (anonymous):

I'm forgetful. SORRY. HAHHAHAHAHA

OpenStudy (aripotta):

for area, you square the amount of times it changed. like, if the scale factor is 3, 3^2 = 9. so the area would be 9 times greater

OpenStudy (anonymous):

Oh. okay.

OpenStudy (aripotta):

if you're doing volume, you cube it

OpenStudy (anonymous):

So the answer would be is 16 times larger than before?

OpenStudy (aripotta):

yea, 16 times the original

OpenStudy (anonymous):

Oh, I GET IT! I THINK. THANKS!

OpenStudy (anonymous):

The perimeter is.

OpenStudy (aripotta):

the perimeter would be 4 times greater

OpenStudy (anonymous):



OpenStudy (anonymous):

Is that a good drawing?

OpenStudy (aripotta):

x is the scale factor. perimeter: x^1, area: x^2, volume: x^3

OpenStudy (anonymous):

thanks. i get it, volume is cubed, area is squared. so the perimeter is .. um..